# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ 3.1 ColME’s Guarantees
[Asadi et al., 2022, Theorem 1] Assume distributions $\{D_{a},\forall a\in\mathcal{A}\}$ are sub-Gaussians with parameter $\sigma^{2}$. For any $\delta\in(0,1)$, and with $\gamma=\frac{\delta}{4|\mathcal{A}|}$, employing ColME, we have:
Turning our attention to the estimation error, it holds:
@@
[Asadi et al., 2022, Theorem 2] Assume distributions $\{D_{a},\forall a\in\mathcal{A}\}$ are sub-Gaussians with parameter $\sigma^{2}$. For any $\delta\in(0,1)$, and with $\gamma=\frac{\delta}{4|\mathcal{A}|}$, employing ColME, we have:
3.2 Extensions
We extend Theorem 1 and 2 to distributions that are not sub-Gaussians as far as the width of the confidence intervals $\beta_{\gamma}(\cdot)$ satisfies the following assumption:
Assumption 1.
There exists a function $\beta_{\gamma}(\cdot)\in o(1)$ such that the true mean belongs to all (nested) intervals of amplitude ±$\pm\beta_{\gamma}(t)$ for $t\in\mathbb{N}$ with confidence $1-2\gamma$, i.e.,
For generic distributions $\{D_{a},a\in\mathcal{A}\}$, it is difficult to find such function $\beta_{\gamma}(\cdot)$, but for sub-Gaussian distributions Eq. (3) satisfies indeed Assumption 1 Maillard [2019].
In Appendix A, we show that Assumption 1 is satisfied when the distributions $\{D_{a},\forall a\in\mathcal{A}\}$ have variance upper-bounded by $\sigma^{2}$, and fourth (central) moment upper-bounded by $\kappa\sigma^{4}$ (when all the variables are identically distributed $\kappa$ can be interpreted as the kurtosis and $\sigma$ as the standard deviation). In particular, in this case $\beta_{\gamma}$, can be selected as follows:
$\beta_{\gamma}(n)=\left(2\frac{\kappa+3\sigma^{4}}{\gamma}\right)^{\frac{1}{4}}\left(\frac{1+\ln^{2}n}{n}\right)^{\frac{1}{4}},$ (7)
If the distributions $\{D_{a},\forall a\in\mathcal{A}\}$ exhibit a larger number of bounded polynomial moments, a more favorable expression for $\beta_{\gamma}(n)$ can be obtained (see Remark 1 in Appendix A).
4 Scalable Algorithms over a Graph $\mathcal{G}$
The main limitations of ColME are acknowledged by Asadi et al. [2022]: each agent needs to perform a number of operations and maintain a memory proportional to $|\mathcal{A}|$, leading to total time and space costs which are $\mathcal{O}(|\mathcal{A}|^{2})$ per time-slot. This is impractical in large-scale systems. Moreover, while the Round-Robin query scheme reduces the communication burden, it entails a significant delay in the estimation ($\zeta_{a}\in\mathcal{O}(|\mathcal{A}|)$). Total space and time computational complexity can be made $\tilde{\mathcal{O}}(1)$ with our scalable approaches: B-ColME (see Section 4.1) and C-ColME (see Section 4.2). Both approaches consider agents $\mathcal{A}$ over a fixed graph $\mathcal{G}(\mathcal{A},\mathcal{E})$ and restrict the communication to pairs of agents adjacent in $\mathcal{G}$. We denote by $\mathcal{CC}_{a}$ the maximal connected component consisting of nodes in $\mathcal{C}_{a}$ to which $a$ belongs, and by $\mathcal{CC}_{a}^{d}$ the agents in $\mathcal{C}_{a}$ which are at most $d$ hops away from $a$. Similarly, let $\mathcal{N}_{a}$ and $\mathcal{N}_{a}^{d}$ denote the set of neighbors of agent $a$ and the set of agents at distance at most $d$ by $a$, respectively.
Each agent $a$ aims to determine which nodes in its neighborhood $\mathcal{N}_{a}$ belong to its similarity class $\mathcal{C}_{a}$. To this purpose, agent $a$ receives at time $t$ from each neighbor $a^{\prime}\in\mathcal{N}_{a}$ an updated version of its local mean estimate. We denote with $\mathcal{C}_{a}^{t}\subseteq\mathcal{N}_{a}$ the set of neighbors that agent $a$ deems to belong to its own similarity class at time $t$. At the beginning, $\mathcal{C}_{a}^{0}=\mathcal{N}_{a}$. Similarly to ColME, at each time $t$, agent $a$ first computes the distance $d_{\gamma}^{t}(a,a^{\prime})$ for every $a^{\prime}\in\mathcal{C}_{a}^{t-1}$ according to (1) and then updates $\mathcal{C}_{a}^{t}$ according to (2). As for ColME, therefore, $\mathcal{C}_{a}^{t}\subseteq\mathcal{C}_{a}^{t-1}$, as soon as $a$ removes $a^{\prime}$ from $\mathcal{C}_{a}^{t}$, it stops communicating with $a^{\prime}$. Then, over time, communications occur over a pruned graph $\mathcal{G}^{t}=(\mathcal{A},\mathcal{E}^{t})$, where $\mathcal{E}^{t}=\{(a,a^{\prime})\in\mathcal{E}:a^{\prime}\in\mathcal{C}_{a}^{t}\}$.
We aim to determine the time needed for all agents in the connected component $\mathcal{CC}_{a}$ to identify their neighbors, i.e., $\mathcal{C}_{a^{\prime}}^{t}=\mathcal{C}_{a^{\prime}}\cap\mathcal{N}_{a^{\prime}}$, $\forall a^{\prime}\in\mathcal{CC}_{a}$. Following the same approach as in [Asadi et al., 2022, Theorem 1] (reported above as Theorem 1), we can prove [^4] Full proofs of this and the following results are in the supplemental material. that:
@@ 4 Scalable Algorithms over a Graph $\mathcal{G}$
For any $\delta\in(0,1)$, employing either B-ColME or C-ColME  we have:
with $\zeta_{a}^{D}=n_{\gamma}^{\star}(a)+1\text{ and }\gamma=\frac{\delta}{4r|\mathcal{CC}_{a}|}$.
Proof.
[…]
When comparing Theorem 3 with Theorem 1, we observe that for large systems, if $r|\mathcal{CC}_{a}|\in\Theta(|\mathcal{A}|)$, $\zeta\approx|\mathcal{A}|+\zeta_{a}^{D}$, showing that, as expected, agents can identify much faster similar agents in their neighborhood than in the whole population $\mathcal{A}$555 For a fairer comparison, we should let ColME query $r$ other agents at each time $t$, where $r$ is the average degree of $\mathcal{G}$. In this case, $\zeta\approx|\mathcal{A}|/r+\zeta_{a}^{D}$ and the conclusion does not change. A more detailed comparison of ColME, B-ColME, and C-ColME is in Sec. 5.
We now present two decentralized algorithms that allow agents to improve the estimates of their true means.
4.1 Message-passing B-ColME
In B-ColME algorithm, each node $a\in\mathcal{A}$, through a continuous exchange of messages with its direct neighbors $\mathcal{C}^{t}_{a}$, can acquire not only their local estimates $\{\bar{x}_{a,a^{\prime}}^{t},a^{\prime}\in\mathcal{C}^{t}_{a}\}$, but also, in an aggregate form, estimates from further away nodes up to distance $d$ in the graph $\mathcal{G}^{t}$ ($d$ is a tunable parameter, the need to introduce it will be explained soon). Indeed, node $a^{\prime}\in\mathcal{C}^{t}_{a}$ serves as a forwarder, providing to its neighbor $a\in\mathcal{N}_{a^{\prime}}$ access to the records it has collected from other neighbors $a^{\prime\prime}\in\mathcal{C}^{t}_{a}\setminus\{a\}$. As far as each agent has correctly identified all similar nodes in its neighborhood, the agent potentially has access to the local estimates of all agents in $\mathcal{CC}_{a}^{d}$.
According to our message-passing scheme, agent $a\in\mathcal{A}$ retrieves at time $t$ a message $M^{t,a^{\prime}\rightarrow a}$ from every of its neighbors $a^{\prime}\in\mathcal{C}^{t}_{a}$ (for a visual representation see Fig. 4 in the Appendix). Message $M^{t,a^{\prime}\rightarrow a}$ is a $d\times 2\,$ table whose element $m^{t,a^{\prime}\rightarrow a}_{h,1}$ contains a sum of samples, while $m^{t,a^{\prime}\rightarrow a}_{h,2}$ indicates the total number of samples contributing to the aforementioned sum. In particular, at time $t$, the first table entries are set as: $m^{t,a^{\prime}\rightarrow a}_{1,1}=\sum_{\tau=1}^{t}x^{\tau}_{a^{\prime}}$ and $m^{t,a^{\prime}\rightarrow a}_{1,2}=t$. The other entries are built through the following recursion:
@@ 4.1 Message-passing B-ColME
for $h\in\{2,\dots,d\}$ and $i\in\{1,2\}$.
If $\mathcal{G}^{t}\cap\mathcal{N}_{a}^{d}$ is a tree, then $m^{t,a^{\prime}\rightarrow a}_{h,1}$ contains the sum of all samples generated within time $t-h+1$ by agents $a^{\prime\prime}\in\mathcal{G}^{t}$ at distance $h-1$ from $a^{\prime}$ and distance $h$ from $a$, while $m^{t,a^{\prime}\rightarrow a}_{h,2}$ contains the corresponding number of samples (the proof is by induction on $h$).
Agent $a$ can estimate its mean as:
@@
and, under the local tree structure assumption, this corresponds to performing an empirical average over all the samples generated by agent $a$ up to $t$ and all agents in $\mathcal{G}^{t}$ at distance $1\leqslant h\leqslant d$ up to $t-h$.
If $\mathcal{G}^{t}\cap\mathcal{N}_{a}^{d}$ is not a tree, then the contribution of the samples collected by a given agent $a^{\prime\prime}$ may be included in messages arriving to $a$ through different neighbors in $\mathcal{C}_{a}^{t}$ and then been erroneously counted multiple times in (9). The parameter $d$ needs to be selected to prevent this eventuality (or make it very unlikely), as we discuss in Sec. 4.3.
As a result of previous arguments, we obtain the following results:
Theorem 4.
. Provided that $\mathcal{CC}^{d}_{a}$ is a tree, for any $\delta\in(0,1)$, employing B-ColME, we have:
$\mathbb{P}\left(\forall t>\tau_{a}^{\prime},\,|\hat{\mu}_{a}^{t}-\mu_{a}|<\varepsilon\right)\geqslant 1-\delta$
$\qquad\text{with}\quad\tau_{a}^{\prime}=\max\left[\zeta_{a}^{D}+d,\frac{\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)}{|\mathcal{CC}^{d}_{a}|}+d\right],$
$\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)\leqslant\left\{\begin{array}[]{ll}\Big{\lceil}-\frac{2\sigma^{2}}{\varepsilon^{2}}\ln\left(\frac{\delta}{4}\Big{(}1-\mathrm{e}^{-\frac{\varepsilon^{2}}{\sigma^{2}}}\Big{)}\right)\Big{\rceil}&\text{sub-Gaussian,}\\ \Big{\lceil}\frac{2(\kappa+3)\sigma^{4}}{\delta\varepsilon^{4}}\Big{\rceil}&\ \text{ otherwise}.\end{array}\right.$
Proof.
The first step consists in evaluating $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)$, defined as the minimum number of i.i.d samples extracted from $D_{a}$, which guarantees that estimate error $|\bar{x}^{t}_{a,a}-\mu_{a}|$ is definitively smaller than $\varepsilon$ with a probability larger por equal than $21-\delta/$ 2. Specifically, we require $\sum_{t=\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)}^{\infty}\mathbb{P}(|\bar{x}^{t}_{a,a}-\mu_{a}|<\varepsilon)>1-\delta/2$. Once $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)$ has been computed, the claim can be easily proved. Indeed under the assumption that $\mathcal{CC}_{a}^{d}$ has been correctly identified a lower bound to the minimum time at which $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)$ i.i.d samples extracted from $D_{a}$ are available at agent $a$ is given by $\frac{\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)}{|\mathcal{CC}_{a}^{d}}|+d$. At last observe that for $t\geqslant\zeta^{D}_{a}+d$, $\mathcal{CC}_{a}^{d}$ has been correctly identified with probability larger than $21-\delta/$ 2, by Theorem 3. The claim follows easily by applying the union bound to the probability of event $\{\exists t\geqslant\tau^{\prime}_{a}:|\mu_{a}^{t}-\mu_{a}|>\varepsilon\}$. ∎
Corollary 5.
$\mathbb{P}\left(\forall t>\tau_{a}^{\prime}\,:\,|\hat{\mu}_{a}^{t}-\mu_{a}|<\varepsilon\right)\geqslant 1-\delta-\delta^{\prime}.$
Now Theorem 23 (in the Appendix) provides upper bounds to $\delta^{\prime}$ when $\mathcal{G}$ is a random regular graph. In particular, as long we set $d$ as in Proposition 8, $\delta^{\prime}$ converges to 00 as the number of agents $|\mathcal{A}|$ increases. Our proofs can readily be adapted to ColME, enabling to substitute $n_{\frac{\delta}{2}}^{\star}(\varepsilon)$ in Theorem 2 with the smaller term $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)$.
@@
4.2 Consensus-based C-ColME
In this section, we present a second possible collaborative approach for estimating the mean, see Algorithm 2. It takes inspiration from consensus algorithms in dynamic settings, as in [Montijano et al., 2014; Franceschelli and Gasparri, 2019]. The basic idea is that each agent maintains two metrics, namely the empirical average of its local samples $\bar{x}_{a,a}^{t}$, and the “consensus" estimate $\hat{\mu}_{a}^{t}$, which is updated at time $t$ by computing a convex combination of the local empirical average and a weighted sum of the consensus estimates in its (close) neighborhood $\{\hat{\mu}_{a^{\prime}}^{t-1},a^{\prime}\in\mathcal{C}_{a}^{t-1}\cup\{a\}\}$.
The dynamics of all estimates are captured by:
where $(W_{t})_{a,a^{\prime}}=0$ if $a^{\prime}\notin\mathcal{C}_{a}^{t}$ and $\alpha_{t}\in(0,1)$ is the memory parameter. Once the agents have discovered which neighbors belong to the same class at time $\tau$, the matrix $W_{t}$ does not need to change anymore, i.e., $W_{t}=W$ for any $t\geqslant\tau$ with $W_{a,a^{\prime}}>0$ if and only if $a^{\prime}\in\mathcal{C}_{a}\cap\mathcal{N}_{a}$. In order to achieve consensus, the matrix $W$ needs to be doubly stochastic Xiao and Boyd [2004] and we also require it to be symmetric.666 This can be obtained in a distributed way, e.g., setting $\forall a^{\prime}~{}\in~{}\mathcal{C}_{a}^{t}$ $(W_{t})_{a,a^{\prime}}=(\max\{|\mathcal{C}_{a}^{t}|,|\mathcal{C}_{a^{\prime}}^{t}|\}\text{+}1)^{-1};\,(W_{t})_{a,a}=1-\sum\limits_{a^{\prime}\in\mathcal{C}_{a}^{t}}(W_{t})_{a,a^{\prime}}.$ By time $\tau$, the original communication graph is split into $C$ connected components, where component $c$ includes $n_{c}$ agents. By an opportune permutation of the agents, we can write the matrix $W$ as follows
where each matrix $\prescript{}{c}{W}$ is an $n_{c}\times n_{c}$ symmetric stochastic matrix. For $t\geqslant\tau$, the estimates in the different components evolve independently. We can then focus on a given component $c$. All agents in the same component share the same expected value, which we denote by $\mu(c)$. Moreover, let $\prescript{}{c}{\bm{\mu}}=\mu(c)\mathbf{1}_{c}$. We denote by $\prescript{}{c}{\mathbf{x}}^{t}$ and $\prescript{}{c}{\hat{\bm{\mu}}}^{t}$ the $n_{c}$-dimensional vectors containing the samples’ empirical averages and the consensus estimates for the agents in component $c$.
To obtain deeper insights into C-ColME and to support the proof of our $(\epsilon,\delta)$ convergence results, it is beneficial to consider an auxiliary system. In this system, the consensus matrix may be arbitrary up until a given time $\tilde{\tau}$, after which it transitions to $W$.
Algorithm 2 C-ColME over a Time Horizon $H$
  $\mathcal{C}_{a}^{0}\leftarrow\mathcal{N}_{a},\forall a\in\mathcal{A}$
  for time $t$ in $\{1,..,H\}$ do
     In parallel for all nodes $a\in\mathcal{A}$
     Draw $x_{a}^{t}\sim D_{a}$
     $\bar{x}^{t}_{a}\leftarrow\frac{t-1}{t}\bar{x}^{t-1}_{a}+\frac{1}{t}x_{a}^{t}$
     Compute $\beta_{\gamma}(t)$ with Eq. (3) or Eq. (7)
     for neighbor $a^{\prime}$ in $\mathcal{N}_{a}\cap\mathcal{C}_{a}^{t-1}$ do
     end for
  end for

### v4

@@ 4 Scalable Algorithms over a Graph $\mathcal{G}$
[Proof in Appendix B] Considering an arbitrarily chosen agent $a$ in $\mathcal{A}$, for any $\delta\in(0,1)$, employing either B-ColME or C-ColME  we have:
with $\zeta_{a}^{D}=n_{\gamma}^{\star}\left(\frac{\Delta_{a}}{4}\right)+1$, $\Delta_{a}=\underset{a^{\prime}\in\mathcal{A}\setminus\mathcal{C}_{a}}{\min}\Delta_{a,a^{\prime}}$, $\gamma=\frac{\delta}{4r|\mathcal{CC}_{a}|}$. $n^{\star}_{\gamma}(x)$ denotes the minimum number of samples that are needed to ensure $\beta_{\gamma}(n)<x$, i.e., $n^{\star}_{\gamma}(x)=\lceil\beta^{-1}_{\gamma}(x)\rceil$.
This result demonstrates that the time required for all agents $a^{\prime}$ in $\mathcal{CC}_{a}$ (the connected component to which $a$ belongs) to correctly identify their neighbors within the same similarity class $\mathcal{C}_{a}$ is bounded by $n^{\star}_{\gamma}(\frac{\Delta_{a}}{4})+1$. Here, $n^{\star}_{\gamma}(\frac{\Delta_{a}}{4})$ represents the number of samples needed to distinguish (with confidence $1-2\gamma$) the true mean of agent $a$ from that of an agent belonging to the ‘closest’ similarity class (i.e., the one with the closest true mean). The additional $1$ accounts for the unit delay in communicating with the neighbors.
When comparing performance of B-ColME and C-ColME (Theorem 1) with ColME (Asadi et al. 2022) [Theorem 1] (reported in Appendix A as Theorem 7 for completeness), we observe that for large systems, if $r|\mathcal{CC}_{a}|\in\Theta(|\mathcal{A}|)$, $\zeta_{a}\approx|\mathcal{A}|+\zeta_{a}^{D}$, showing that, as expected, agents can identify much faster similar agents in their neighborhood than in the whole population $\mathcal{A}$111 For a fairer comparison, we should let ColME query $r$ other agents at each time $t$, where $r$ is the average degree of $\mathcal{G}$. In this case, $\zeta_{a}\approx|\mathcal{A}|/r+\zeta_{a}^{D}$ and the conclusion does not change. See Sec. 5 for a detailed comparison of ColME, C-ColME, and B-ColME.
4.1 Consensus-based Algorithm: C-ColME
This section introduces the first collaborative mean estimation approach, inspired by consensus algorithms in dynamic settings, as in (Montijano et al. 2014; Franceschelli and Gasparri 2019). The basic idea is that each agent maintains two metrics: the empirical average of its local samples $\bar{x}_{a,a}^{t}$, and the ‘consensus’ estimate $\hat{\mu}_{a}^{t}$. The consensus variable is updated at time $t$ by computing a convex combination of the local empirical average $\bar{x}^{t}_{a,a}$ and a weighted sum of the consensus estimates in its (close) neighborhood $\{\hat{\mu}_{a^{\prime}}^{t-1},a^{\prime}\in\mathcal{C}_{a}^{t-1}\cup\{a\}\}$, see Algorithm 1.
The dynamics of all estimates are captured by:
where $(W_{t})_{a,a^{\prime}}=0$ if $a^{\prime}\notin\mathcal{C}_{a}^{t}$ and $\alpha_{t}\in(0,1)$ is the memory parameter. Once the agents cease pruning their neighbors, say at time $\tau$, the matrix $W_{t}$ does not need to change anymore, i.e., $W_{t}=W$ for any $t\geqslant\tau$ with $W_{a,a^{\prime}}>0$ if and only if $a^{\prime}\in\mathcal{C}_{a}\cap\mathcal{N}_{a}$. In order to achieve consensus, the matrix $W$ needs to be doubly stochastic (Xiao and Boyd 2004) and we also require it to be symmetric. By time $\tau$, the original communication graph is split into $C$ connected components, where component $c$ includes $n_{c}$ agents. By an opportune permutation of the agents, we can write the matrix $W$ as follows
where each matrix $\prescript{}{c}{W}$ is an $n_{c}\times n_{c}$ symmetric stochastic matrix. For $t\geqslant\tau$, the estimates in the different components evolve independently. We can then focus on a given component $c$. All agents in the same component share the same expected value, which we denote by $\mu(c)$. Moreover, let $\prescript{}{c}{\bm{\mu}}=\mu(c)\mathbf{1}_{c}$. We denote by $\prescript{}{c}{\mathbf{x}}^{t}$ and $\prescript{}{c}{\hat{\bm{\mu}}}^{t}$ the $n_{c}$-dimensional vectors containing the samples’ empirical averages and the consensus estimates for the agents in component $c$ and by $\lambda_{2,c}$ the second largest module of the eigenvalues of $\prescript{}{c}{W}$.
Note that the actual evolution of $W_{t}$ is challenging to characterize due to topology changes during the graph pruning phase. However, our main results (Theorems 2 and 3) remain applicable to any system where the sequence of matrices $W_{t}$ for $t\leq\tau$ is arbitrarily set.
Algorithm 1 C-ColME over a Time Horizon $H$
  $\mathcal{C}_{a}^{0}\leftarrow\mathcal{N}_{a},\forall a\in\mathcal{A}$
  for time $t$ in $\{1,..,H\}$ do
     In parallel for all nodes $a\in\mathcal{A}$
     Draw $x_{a}^{t}\sim D_{a}$
     $\bar{x}^{t}_{a}\leftarrow\frac{t-1}{t}\bar{x}^{t-1}_{a}+\frac{1}{t}x_{a}^{t}$
     Compute $\beta_{\gamma}(t)$ with Eq. (3) or Eq. (5)
     for neighbor $a^{\prime}$ in $\mathcal{N}_{a}\cap\mathcal{C}_{a}^{t-1}$ do
     end for
  end for
@@ 4.1 Consensus-based Algorithm: C-ColME
[Proof in Appendix F] Consider a system which evolves according to (7) with $W_{t}=W$ in (8), for $t\geqslant\tau$. Let $\prescript{}{c}{P}=1/n_{c}\mathbf{1}_{c}\mathbf{1}_{c}^{\intercal}$. For $\alpha_{t}=\frac{t}{t+1}$, it holds:
$+\mathcal{O}\left(\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]\left(\frac{1+\ln t}{1+t}\right)^{2}\right).$
The theorem shows that the error, quantified through the fourth moment, can be decomposed into three terms decreasing over time. The first term depends on the estimates’ error at time $\tau$. The second term captures the effect of the consensus averaging, i.e., how effective is the algorithm in bringing the local estimates $\prescript{}{c}{\mathbf{x}}$ close to their empirical value $\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}$ (for example it is minimized if $\lambda_{2}=0$, which corresponds to $\prescript{}{c}{W}=\prescript{}{c}{P}$, the ideal choice for the matrix $\prescript{}{c}{W}$). Finally, the third term represents the minimum possible error, which would be obtained by averaging the estimates of all agents in the component using the matrix $\prescript{}{c}{P}$.
Theorem 3 shows that C-ColME achieves a speedup proportional to the size of the connected component $|\mathcal{CC}_{a}|$.
@@
[Proof in Appendix F] Consider a graph component $c$ and pick uniformly at random an agent $a$ in $c$. Let $g(x)\coloneqq x\ln^{2}(ex)$ and $\alpha_{t}=\frac{t}{t+1}$. Under BFMD, it holds:
$\mathbb{P}\left(\forall t>\tau^{C}_{a},\,|\hat{\mu}_{a}^{t}-\mu_{a}|<\epsilon\right)\geqslant 1-\delta$
where $\tau_{a}^{C}=\max\left\{\zeta_{a}^{D},g\!\left(C\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{|\mathcal{CC}_{a}|\epsilon^{4}\delta}\right)\in\tilde{\mathcal{O}}\left(\frac{\tilde{n}_{\frac{\delta}{2}}(\varepsilon)}{|\mathcal{CC}_{a}|}\right)\right\}$ and $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)=\Big{\lceil}\frac{2(\kappa+3)\sigma^{4}}{\delta\varepsilon^{4}}\Big{\rceil}$.
The theorem shows that the time to reach an $\epsilon$-accurate estimate with high probability is the maximum of the time for the agents in $\mathcal{CC}_{a}$ to identify their neighbors in the same similarity class and the time required for those agents to obtain and $\epsilon$-accurate estimate if they could share their own samples. Indeed, we observe that $n_{\delta/2}^{\star}(\epsilon)$ is the number of samples sufficient to ensure that $\mathbb{P}(|\hat{\mu}_{a}-\mu_{a}|>\epsilon)<\delta/{2}$ (see details in Appendix E) and that the nodes in $\mathcal{CC}_{a}$ collectively gather this number of samples by time $t=\Big{\lceil}\frac{n_{\delta/2}^{\star}(\varepsilon)}{|\mathcal{CC}_{a}|}\Big{\rceil}$.
Appendix F also presents convergence results for the case $\alpha_{t}=\alpha$, but they do not enjoy the same speedup factor.
4.2 Message-passing Algorithm: B-ColME
In B-ColME, each node $a\in\mathcal{A}$ continuously exchanges messages with its direct neighbors $a^{\prime}\in\mathcal{C}^{t}_{a}$. This enables node $a$ to acquire not only the neighbor’s local estimates $\{\bar{x}_{a,a^{\prime}}^{t},a^{\prime}\in\mathcal{C}^{t}_{a}\}$, but also aggregated estimates from nodes up to a distance $d$ in the graph $\mathcal{G}^{t}$ (where $d$ is a tunable parameter). Indeed, each neighbor $a^{\prime}$ acts as a forwarder, granting node $a$ access to the records from its own neighbors $a^{\prime\prime}\in\mathcal{C}^{t}_{a^{\prime}}\setminus\{a\}$. Provided each agent correctly identifies all similar nodes in its neighborhood, agent $a$ can potentially access the (delayed) local estimates of all agents in $\mathcal{CC}_{a}^{d}$.
In our message-passing scheme, at time $t$, agent $a\in\mathcal{A}$ receives a message $M^{t,a^{\prime}\rightarrow a}$ from all neighbors $a^{\prime}\in\mathcal{C}^{t}_{a}$. The message $M^{t,a^{\prime}\rightarrow a}$ is a $d\times 2\,$ table whose elements $m^{t,a^{\prime}\rightarrow a}_{h,1}$ contain a sum of samples, while $m^{t,a^{\prime}\rightarrow a}_{h,2}$ indicates the number of samples contributing to this sum. In particular, at each time $t$, the first row of the table is set as: $m^{t,a^{\prime}\rightarrow a}_{1,1}=\sum_{\tau=1}^{t}x^{\tau}_{a^{\prime}}$ and $m^{t,a^{\prime}\rightarrow a}_{1,2}=t$, i.e., the immediate neighbors’ sum of local samples. The remaining entries are computed through the following recursion:
@@ 4.2 Message-passing Algorithm: B-ColME
for $h\in\{2,\dots,d\}$ and $i\in\{1,2\}$. This captures information extending beyond immediate neighbors. For additional details on B-ColME see Algorithm 2 and Fig. 5 in Appendix D.
If $\mathcal{G}^{t}\cap\mathcal{N}_{a}^{d}$ is a tree, then $m^{t,a^{\prime}\rightarrow a}_{h,1}$ contains the sum of all samples generated within time $t-h+1$ by agents $a^{\prime\prime}\in\mathcal{G}^{t}$ at distance $h-1$ from $a^{\prime}$ and distance $h$ from $a$, while $m^{t,a^{\prime}\rightarrow a}_{h,2}$ contains the corresponding number of samples (the proof is by induction on $h$). Agent $a$ can estimate its mean as:
@@
Under the local tree structure assumption, this corresponds to performing an empirical average over all the samples generated by all agents in $\mathcal{G}^{t}$ at distance $0\leqslant h\leqslant d$ from $a$ up to time $t-h$. If $\mathcal{G}^{t}\cap\mathcal{N}_{a}^{d}$ is not a tree, samples collected by a given agent $a^{\prime\prime}$ may be included in messages received by $a$ through different parallel paths (from $a$ to $a^{\prime\prime}$). As a result, these samples are erroneously counted multiple times in (9). The parameter $d$ must be chosen to prevent this issue with high probability, as discussed in Sec. 4.3.
@@
Theorem 4 presents the $(\epsilon,\delta)$ convergence result for B-ColME, which enjoys a speedup proportional to $|\mathcal{CC}^{d}_{a}|$.